Week of 181203

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Miro (WLCG, chair), Kate (WLCG), Maarten (WLCG, ALICE), Borja (monit), Robert (LHCb), Christoph (CMS), Andrei (DB), Rémy (Storage), Gavin (Computing), Alberto (monitoring)
  • remote: Marcelo (CNAF), David B (IN2P3), Xavier (KIT), Christian (NDGF), Andrew (NL-T1), Pepe (PIC), John (RAL), Di (TRIUMF)

Experiments round table:

  • CMS reports ( raw view) -
    • Heavy Ion left us (as expected) with a back log of 'prompt' reconstruction and data archiving (mainly to FNAL)
      • RAW data saving at CERN kept up without major delay
      • Reconstruction and archiving expected to last until Xmas or beyond
    • No major trouble

  • ALICE -
    • Successful end of the Run 2 data taking!
      • Reconstruction and reprocessing for many months to come
      • Plus lots of analysis and MC simulation as usual
    • Xavier: due to high load from ALICE on the tape system, the SAM tests for the tape SE are failing
    • Maarten: we will correct the A/R numbers as needed (done after the meeting)

  • LHCb reports ( raw view) -
    • Activity
      • Data reconstruction for 2018 data
      • User and MC jobs
    • Site Issues
      • SARA: Ticket open concerning data transfers problems (GGUS:138472) site waiting on CERN input. CERN network team now involved in IPv6 troubleshooting.

Sites / Services round table:

  • ASGC: nc
  • BNL: nc
  • CNAF:
    • During the weekend there was an issue with the tape migration. The buffer got full and all the operations after that failed till the buffer was empty. This affected ATLAS and ALICE. Now all seems good.
    • CMS xRootd loop fixed. xrootd redirector Fallback
  • EGI: nc
  • FNAL: nc
  • IN2P3: In downtime tomorrow (4th December) for the whole day as announced in WLCGOpsMeetingWeek181126
  • JINR: Minor LAN problem. Under investigation.
  • KISTI: Next week, KISTI will have a scheduled downtime from 9th December 23:00 CET (10th December 07:00 KST) to 14th December 10:00 CET (14th December 18:00 KST) due to facility migration. Our whole system will be shutdown, moved to new building and then recovered during the downtime.
  • KIT: NTA
  • NDGF: NTR
  • NL-T1: Ongoing webdav connection issues to the nikhef DPM server, see (GGUS:138476).
  • NRC-KI: nc
  • OSG: nc
  • PIC: Scheduled downtime on 11th of December. dCache will be down.
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services:
  • CERN storage services:
    • EOS : Point5 to EOSCMS: Poor performance overnight, found that data distribution on the cluster is less than ideal, focusing new writes on a limited subset of machines. We are working actively to improve the situation.
    • CASTOR : Reminder for Alice : Please send data to tape.

  • CERN databases:
  • GGUS: NTR
  • Monitoring:
    • Issue with SAM3 CMS availability computations has been understood and the availabilties are being computed right now. More details will be given on the open tickets.
  • MW Officer: nc
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2018-12-04 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback